OpenStudy (anonymous):

A shopping bag contains 6 orange juice boxes and 4 grape juice boxes. Without looking, April selects one juice box, hands it to her brother, and then takes a second juice box for herself. What is the probability that April gets an orange juice box and her brother gets a grape juice box?

jigglypuff314 (jigglypuff314):

I'm not good at probability, so I might be wrong but I think that it would be (probability of first round) * (probability of second round) probability of first round = grape out of total = 4/10 probability of second round = orange out of remaining = 6/9 so (4/10) * (6/9) = ?